What is the distribution of stresses in an artery that has internal stresses such that (a) α = 180°; (b) α = 150°?At what internal pressure will the stress outside and inside the wall become the same? Assume (i) that the stress from the pressure decays linearly to zero at the external surface, and (ii) a linear elastic behavior with E = 400 MPa. Given: ID = 15 mm; OD = 22 mm.


1
Expert's answer
2021-08-19T03:21:01-0400

p = internal pressure

"\sigma= \frac{P}{A}"

"E\times e= \frac{P}{A}"


"P =E\times e\times A = 400 \times 1\times \frac{\pi}{4}(22^2-15^2) =81.36kN"